OpenStudy (andrewkaiser333):

Solve for X. (2x+5)(6x-1)=120

OpenStudy (taylor0402):

(2x + 5)(6x - 1) = 120 12x^2 - 2x + 30x - 5 = 120 12x^2 + 28x - 125 = 0 x = [-28 +/- sqrt(784 + 4*12*125)]/24 x = [-28 +/- sqrt(6784)]/24 x = [-28 +/- 8sqrt(106)]/24 x = [-7 +/- 2sqrt(106)]/6 Though they are ugly, these are your two solutions to the original equation.

OpenStudy (andrewkaiser333):

can u do it take a pic of it and attach it so i see what it is meaning i dont understand this

OpenStudy (taylor0402):

no because im on my computer so, go on google and search tiger algebra and plug it in on there it should give you a better explanation

OpenStudy (andrewkaiser333):

this makes no sense

OpenStudy (wolf1728):

Andrew, do you know how to multiply the equation? (2x+5)(6x-1)=120

OpenStudy (andrewkaiser333):

no its been too long

OpenStudy (wolf1728):

That would be the first step. Taylor0402 already multiplied it for you and it is 12x^2 - 2x + 30x - 5 = 120

OpenStudy (wolf1728):

Andrew if it has been a long time since you have done algebra, are you even ready to be doing this type of math?

OpenStudy (andrewkaiser333):

i am doin this to help my gf. My memory loss is whats making it harder to remember its been 3 years since i did this that i can remember and am wanting to learn it again so i can help her with her homework and get better at this again. i had straight a's when i had done it but yeah.

OpenStudy (wolf1728):

Okay, the next step is to get the equation 12x^2 - 2x + 30x - 5 = 120 and add the terms so you can have an equation with just 3 terms. Can you do that?

OpenStudy (andrewkaiser333):

other subjects from this i know still because i use it everyday

OpenStudy (wolf1728):

Can you do what I asked you in the posting above?

OpenStudy (andrewkaiser333):

12x^2-2x+30x-125=0?

OpenStudy (wolf1728):

Okay but you have to combine the -2x with +30x

OpenStudy (andrewkaiser333):

oh yeah 12x^2+28-125=0

OpenStudy (wolf1728):

Now you solve the equation using the Quadratic Formula

OpenStudy (wolf1728):

The "a" term is 12 The "b" term is 28 The "c" term is -125

OpenStudy (andrewkaiser333):

\[\frac{ -28+/-\sqrt{28^2-4(12x-125)} }{ 2a }\]

OpenStudy (andrewkaiser333):

yes?

OpenStudy (wolf1728):

That should be 28^2 -4*12*-125 and 2a should be 2* 12 or 24

OpenStudy (andrewkaiser333):

use equation button on here please so i can understand

OpenStudy (wolf1728):

\[4 a c = 4*12*-125\]

OpenStudy (wolf1728):

It is NOT 4 *(12 -125)

OpenStudy (andrewkaiser333):

ok

OpenStudy (wolf1728):

and 2a = 2 * 12 or 24

OpenStudy (andrewkaiser333):

can u make the whole thing i did with the fixes so i see how it should be?

OpenStudy (wolf1728):

\[x = (-28 ± \sqrt{28² - 4*12*-125} ) ÷ 2*12\]

OpenStudy (wolf1728):

\[x = (-28 ±\sqrt{784 +6,000} ) ÷ 24\]

OpenStudy (andrewkaiser333):

ok

OpenStudy (wolf1728):

Okay then

OpenStudy (andrewkaiser333):

what next

OpenStudy (wolf1728):

\[x = (-28 ±\sqrt{6,784} ) ÷ 24\]

OpenStudy (wolf1728):

Taylor 0402 pretty much did what I just did so look for that solution almost at the top of this posting.

OpenStudy (andrewkaiser333):

thanks

OpenStudy (andrewkaiser333):

now i understand